OpenStudy (anonymous):

Given f(x) = x + 2 and g(x) = x2 – 4, find the function (g/f)(x).

OpenStudy (anonymous):

Recall that : \[\left(\frac gf\right)(x)=\frac{g(x)}{f(x)}\]

OpenStudy (dan815):

^yes do that and simplify

OpenStudy (jhannybean):

i would first simplify (x^(2) -4) This is a difference in squares where (a-b)(a+b) how would you break this down?

OpenStudy (anonymous):

your excluded value u wanted to know was 

OpenStudy (anonymous):

just caz my hand writing is SHOCKING

OpenStudy (jhannybean):

....:| i ant even read it. I've learned to read @dan815 's...but thats..... x_x.''

OpenStudy (anonymous):

it says x is any real number excluding -2!

OpenStudy (anonymous):

haha

OpenStudy (jhannybean):

nooo don't give the answer! >:(

OpenStudy (anonymous):

im using my laptop. i got no mouse bro

OpenStudy (anonymous):

did u think i was a ret?

OpenStudy (dan815):

lolol close enuff

OpenStudy (anonymous):

oops I accidentally asked the last question guys.

OpenStudy (anonymous):

close enuff? COME ON!

OpenStudy (dan815):

but what i wanna know is how did you know it was Denis the menace

OpenStudy (anonymous):

i code lines as a living

OpenStudy (jhannybean):

...sure.